# Writing – review & editing, Writing – original draft, Visualization, Data curation ,Investigation, Formal analysis, Validation, Software, Methodology, Conceptualization

### v2

@@ 4.5 Attack Configuration
The transferability-based attacks require the attacker to train a surrogate model (illustrated in Fig. 1 (d, e)). In our evaluation, all surrogate models were neural networks, and we used an embedding layer to try to maintain the consistency of the crafted samples, following the approach employed in other studies [36, 21]. All surrogate models were comprised of two components: a sub-model for embedding and a sub-model for classification, both of which were fine-tuned for optimized performance. The embedding sub-model had an input size matching the number of features in the input sample, followed by a dense layer with 256 neurons and ReLU and PReLU activations for the Hate and VideoTQ datasets respectively. For the ICU dataset, the embedding sub-model had three dense layers with 256, 128, and 64 neurons and ReLU activation for all layers. The output size of all embedding sub-models was set at 16, i.e., the embedding size. The classification sub-model had a dense layer with 16 neurons (matching the embedding size) with ReLU activation for the Hate and ICU datasets and PReLU activation for the VideoTQ datasets. This was followed by a dropout layer with a dropout rate of 0.1. All surrogate models were trained using the binary cross-entropy loss function [32] and the Adam optimizer [28] with a learning rate of 0.2 for the Hate and VideoTQ datasets and 0.1 for the ICU dataset.
After training, the attacker performs the adversarial attack against the surrogate model by perturbing a group of features in each iteration. Each transferability-based attack employs a different feature selection technique to determine which features to perturb (see Appendix A).
For importance-based selection technique, the $k$ most important features are selected, along with the $n$ features most correlated with them—resulting in a total of $k\times n$ additional perturbed features per iteration. In each subsequent iteration, the next $k$ most important features and their $n$ most correlated features are added to the set of perturbed features. This process continues until an adversarial sample is found or no additional features remain for selection. Feature importance is determined using XGBoost, gradient boosting, LightGBM, and random forest models, each trained separately for this purpose. For random-based selection attacks, $k$ features are randomly chosen in each iteration, along with their $n$ most correlated features.
For all attacks, $k$ was set at two and $n$ was set at one (i.e., one correlated feature per selected feature). The correlated features were identified using the Pearson correlation coefficient [49].
@@ 4.6 Regression Models
Of the datasets used in our evaluation, the ICU dataset is the only one containing dependent features whose dependencies are not well-defined. Specifically, these dependent features are: ‘apache_3j_bodysystem,’ ‘apache_3j_diagnosis,’ ‘d1_mbp_invasive_max,’ and ‘d1_mbp_invasive_min.’
The regression models were trained on the dataset used to train the surrogate models, using 200 estimators, a max depth of six, and a learning rate of 0.1 for the ‘apache_3j_bodysystem’ and ‘apache_3j_diagnosis’ features, and 0.01 for the ‘d1_mbp_invasive_max’ and ‘d1_mbp_invasive_min’ features.
@@ 4.8 Evaluation Setup
To ensure a robust evaluation, we created the attack set by filtering the test set to include only samples correctly classified by both the target and surrogate models, excluding those already misclassified. This filtering retained $85\%$, $69\%$, and $91\%$ of samples from the Hate, ICU, and VideoTQ datasets, respectively. Next, we randomly selected a balanced number of samples from each class. As a result, the attack sets contained 182, 1,000, and 1,000 samples for the Hate, ICU, and VideoTQ datasets, respectively. Our analysis of adversarial samples (using $L_{p}$ norms or quality assessment), considered only samples that successfully fooled the target model (or both the surrogate and target models in the case of transferability-based attacks).
@@ 4.9 Evaluation Metrics
The main objective of our evaluation is to analyze the characteristics of and differences between black-box query- and transferability-based adversarial attacks with respect to three key factors: the attacker’s risk, the attacker’s effort, and the attack quality (illustrated in Fig. 1(f)). To perform a thorough analysis, we used the evaluation metrics commonly used in the literature and the metrics proposed in this work.
To assess the attacker’s risk we used three metrics (illustrated in Fig. 1(f-1)): i) the percentage of samples that successfully misled the model (attack success rate); ii) the average number of modified features ($L_{0}$) which might influence the attack’s ability to be successfully performed in real-life tabular data domains (due to the possible cost of changing the features), and the average Euclidean distance between the original and adversarial samples ($L_{2}$); and iii) the average number of queries required.
To assess the attacker’s effort we used three metrics (illustrated in Fig. 1(f-2)): i) the time it took to craft an adversarial sample; ii) the number of data samples needed to execute the attack; and iii) the memory resources and computing capabilities required to craft the adversarial sample and perform the optimization process.
The attack quality is reflected in the crafted adversarial sample, which were assessed using three metrics (illustrated in Fig. 1(f-3)): i) the anomaly detection rate of adversarial samples compared to benign samples, using IF and AE models, where a higher rate indicates less consistency and coherence; ii) the percentage of samples with at least one feature with an anomalous SHAP value (Importance-Based Anomaly Detection Rate); and iii) the average anomalous SHAP features per sample.
In all three metrics, a lower number is better as higher values indicates an abnormal influence on the model’s decision-making process. Each experiment was conducted separately for each target model, as SHAP values vary across models.
@@ 5.1 Models’ Performance
Table 1 summarizes the performance of all target and surrogate models examined, presenting the accuracy, F1 score, precision, and recall values obtained by each model on the test set. As can be seen, all models were high-performing. The target models for the Hate dataset obtained accuracy values ranging from [$88\%-91\%$], F1 scores ranging from [$87\%-92\%$], precision values ranging from [$89\%-94\%$], and recall values ranging from [$82\%-92\%$]. The target models for the ICU dataset obtained accuracy values ranging from [$77\%-80\%$], F1 scores ranging from [$75\%-81\%$], precision values ranging from [$81\%-84\%$], and recall values ranging from [$69\%$-$77\%$]. The target models for the VideoTQ dataset obtained accuracy values ranging from [$98\%-99\%$], F1 scores ranging from [$99\%-100\%$], precision values ranging from [$91\%-97\%$], and recall values ranging from [$96\%$-$99\%$].
@@ 5.2 Attacker’s Risk
Tables 2 and 3 and Fig. 2 present the results for the attacker’s risk metrics (see Section 4.9) for the query- and transferability-based attacks performed against different target models on each dataset.
Table 2 presents details on the following attack success rates: Target Success Rate (SR): The percentage of adversarial samples that successfully mislead the target model; Surrogate Success Rate (Surrogate SR): The percentage of adversarial samples that successfully mislead the surrogate model; Transfer Success Rate (Transfer SR): The percentage of adversarial samples that successfully transfer from the surrogate model to the target model, calculated as the ratio of the Surrogate SR. Overall Success Rate (Overall SR): The percentage of adversarial samples that successfully mislead both the surrogate and target models, computed as the ratio of the entire attack set.
@@
[…]
Table 3 presents the average number of queries required to generate adversarial samples for query-based attacks. As can be seen, the examined query-based attacks require a significant number of queries, ranging from hundreds to hundreds of thousands. In contrast, transferability-based attacks require just a single query for all datasets.
@@
Fig. 2 illustrates the distribution of both the number of modified features ($L_{0}$ ) and the Euclidean distance ($L_{2}$) between the adversarial attacks and the original samples.
@@ 5.3 Attacker’s Effort
To assess the attacker’s effort, we compared the performance of the query- and transferability-based attacks in terms of the computational time, the required amount of data, and the computational resources needed(see Section 4.9).
With respect to the computational time, the attacker’s objective is to minimize the amount of time it takes to craft an adversarial sample. In query-based attacks, the computational time includes: i) the time it takes to query the target model (denoted as $\alpha$); and ii) the time it takes to perform a single optimization iteration (denoted as $\beta$). The total amount of time it takes to craft an adversarial sample is calculated in Eq. 2:
@@
where $x$ is a data sample, $M$ is the target model, $n$ is the number of queries required to craft a successful adversarial sample, and $t$ is the time it takes to perform a single query. When considering solely the query-related aspects of the required time, query-based attacks are at a disadvantage compared to transferability-based attacks, since many more queries to the target model are required (see Section 5.2). However, in transferability-based attacks, there is another aspect to consider – the surrogate model’s training and querying time. The total amount of time it takes to craft an adversarial sample is calculated in Eq. 3, where we assume that the time it takes to perform a query to the target model is identical to the time required to perform a query to the surrogate model.
@@
In our evaluation, we found that the total amount of time it took to produce a single adversarial sample in query-based attacks was, on average, $518.4$ seconds. However, in transferability-based attacks, it was $\sim$243.5673243.5673243.5673243.5673 seconds: $\sim$240.49240.49240.49240.49 seconds to train the surrogate and feature importance models, $0.0773$ seconds to generate the sample w the surrogate model, and $\sim$3333 seconds to query the target model. Based on this, we can conclude that although transferability-based attacks require training additional models, the total computational time required to craft a single adversarial sample is shorter than in query-based attacks.
@@
In conclusion, while query-based attacks require less data, transferability-based attacks require less computational time and resources. However, they are limited by lower success rates compared to query-based attacks, as discussed in Section 5.2.
@@ 5.4 Attack Quality
In our evaluation, we assessed query- and transferability-based attacks in terms of the quality of the adversarial samples produced. This quality can be examined by considering both the adversarial sample’s coherence and its impact on the target model’s decision-making process. To measure the coherence of adversarial samples, we propose using metrics based on AE and IF models. To evaluate the impact on the target model’s decision-making process, we introduced two new metrics derived from the SHAP framework (see Section 4.9).
Figure 3: Attack quality evaluated based on the anomaly detection rate by AE and IF models across datasets, computed separately for each class. The FPR on all datasets for both classes is less than 0.04.
@@
Feature Space Coherency.  Fig. 3 presents the anomaly detection rates derived from the values obtained using the IF model and the reconstruction errors from the AE for adversarial samples for each evaluated attack, target class, and dataset. Each anomaly detection rate represents the percentage of adversarial samples identified as anomalies. For the AE, a sample is considered anomalous if its reconstruction error exceeds a predefined threshold (see  Eq. 1. The false positive rate (FPR) for benign samples, based on the specified threshold, is 0.020 and 0.011 for class 0 and class 1 on the Hate dataset, 0.038 and 0.038 on the ICU dataset, and 0.014 and 0.002 on the VideoTQ dataset. The same FPR was applied to the IF model for comparison.
On the Hate dataset, the boundary attack led to a $100\%$ detection rate for adversarial samples for both target classes by both the AE and IF models. For the HopSkipJump attack, a $100\%$ anomaly detection was obtained for adversarial samples in class 0 by both the AE and IF models when targeting the GB and LGB models; however, no anomalies were detected for adversarial samples in class 1 across all target models. This discrepancy may arise from HopSkipJump’s initial adversarial sample selection and the optimization process, i.e., selecting a sample from the target class that is closest to the original sample and increasing proximity to the original sample using gradient estimation, whereas the boundary attack uses a random initial adversarial sample from the target class and does not employ gradient estimation (see Appendix A). In the case of the transferability-based attacks, only the AE successfully detected anomalies, achieving a $100\%$ anomaly detection rate for most attacks for both target classes.
On the ICU dataset, the boundary attack demonstrated similar performance trends as observed for the Hate dataset, however, the HopSkipJump attack resulted in fewer detected anomalies for both target classes across both the AE and IF models. The transferability-based attacks also led to significantly fewer anomalies, with the exception of the LGB importance-based (LGB imp.) and random-based selection attacks. These two methods resulted in 60–90% of samples from both target classes being detected as anomalies by the AE model.
On the VideoTQ dataset, all attacks resulted in high anomaly detection rates (85–100%) by the AE model. Except for the boundary attack, which achieved a 0% anomaly detection rate for class 1 by both the AE and IF models.
@@
As shown in Fig. 3, the two query-based attacks demonstrate distinct behaviors in terms of the models’ anomaly detection rates. Most adversarial samples generated by the boundary attack were detected as anomalous by both models. In contrast, adversarial samples from the HopSkipJump attack, similar to those from transferability-based attacks, were consistently detected as anomalous only by the AE model.
@@
Notably, despite the HopSkipJump and transferability-based attacks producing samples with lower $L_{2}$ distances (as detailed in Section 5.2, these samples were still identifiable as anomalous. This suggests that even subtle perturbations introduced by these attacks can still be identified as anomalous.
@@
Figs. 4, 5 and 6 presents the distribution of SHAP values for adversarial versus benign samples across different attacks and datasets for each target model. To simplify the presentation, we focus on the top-four most important features, selected based on their average SHAP values computed on the benign training dataset.
As shown in the figures, for most target models and datasets, there is at least one feature where the SHAP value distribution across attacks varies significantly. Analyzing the SHAP-based feature importance of a queried sample can therefore be an effective technique for identifying attacks.
Table 4 presents the results of our evaluation of attack quality using SHAP-based metrics. For each attack and target model, we report the Importance-Based Anomaly Detection Rate (percentage of samples with at least one anomalous feature based on SHAP values) and the Average Anomalous SHAP Features per Sample (as described in Section 4.9). Lower values indicate that there is less impact on the model’s decision-making process, meaning that adversarial samples were more consistent with benign samples in terms of feature importance (i.e., the model relies on similar features for prediction).
@@
On the ICU dataset, both query- and transferability-based attacks generally produced less anomalous SHAP values for most target models (LGB, XGB, and RF), with importance-based anomaly detection rates typically below 7% and very low values of average anomalous SHAP features per sample ranging from 0.0 to 0.1. However, when targeting the GB model, transferability-based attacks resulted in notably lower importance-based anomaly detection rates (1.8% to 10.8%) compared to query-based attacks (29.0% and 19.4% for the boundary and HopSkipJump attacks, respectively). In this case, the average number of anomalous SHAP features per sample was slightly higher for query-based attacks (0.3 to 0.4) than for transferability-based attacks (0.0 to 0.1), though still relatively low overall.

### v3

@@ 4.4 Attack Configuration
The transferability-based attacks require the attacker to train a surrogate model in order to generate adversarial samples that are transferred to the target black-box model (illustrated in Fig.˜1). In our evaluation, all surrogate models were neural networks in which an embedding layer was used to try to maintain the consistency of the crafted samples, following the established approaches [8, 9]. The surrogate models were comprised of two components: a sub-model for embedding and a sub-model for classification, both of which were fine-tuned for optimized performance.
@@
The embedding sub-model had an input size matching the number of features in the input sample, followed by a dense layer with 256 neurons and ReLU and PReLU activations for the Hate and VideoTQ datasets respectively. For the ICU dataset, the embedding sub-model had three dense layers with 256, 128, and 64 neurons and ReLU activation for all layers. The output size of all embedding sub-models was set at 16, i.e., the embedding size. The classification sub-model had a dense layer with 16 neurons (matching the embedding size), with ReLU activation for the Hate and ICU datasets and PReLU activation for the VideoTQ datasets, followed by a dropout layer with a dropout rate of 0.1. All surrogate models were trained using the binary cross-entropy loss function [55] and the Adam optimizer [56] with a learning rate of 0.2 for the Hate and VideoTQ datasets and 0.1 for the ICU dataset. Training was conducted for a maximum of 70 epochs, with early stopping applied based on the validation loss (patience = 2 epochs). The model parameters from the epoch yielding the lowest validation loss were retained.
After training, the attacker performs the adversarial attack against the surrogate model using specific feature selection techniques. We implemented four attacks that select features for perturbation based on their importance scores, differing only in the model used to compute the feature importance (via SHAP values): XGBoost, gradient boosting, LightGBM, or random forest, each trained separately for this purpose. Additionally, we implemented a random-based selection attack that selects features uniformly at random. In all attacks, $k=2$ features are selected per iteration along with $n=1$ correlated features (identified using the Pearson correlation coefficient [57]), continuing until an adversarial sample is found or the maximum allowed $L_{0}$ perturbation $\lambda$ is reached. In our implementation $\lambda$ is set to be the number of editable features.
The adversarial objective underlying these attacks is to generate adversarial samples that mislead the surrogate model while preserving similarity to the original input in the latent space. This is achieved by minimizing the adversarial loss function, defined as:
$\begin{split}L_{adv}(x_{adv},x,y)=-\mathrm{BCE}(M^{\prime}(x_{adv}),y)\\ +\alpha\left\|\phi(x_{adv})-\phi(x)\right\|_{2}\end{split}$ (2)
where $y$ is the true label of $x$, $\mathrm{BCE}$ denotes the binary cross-entropy loss, $M^{\prime}$ is the surrogate model, $\phi(\cdot)$ is the embedding sub-model that outputs the latent representation (embedding), and $\alpha=1$ is the regularization parameter. The first term promotes misclassification by maximizing the classification error of $x_{adv}$, whereas the second term ($L_{2}$ regularization) seeks to preserve similarity by minimizing the Euclidean distance between the latent representations of $x_{adv}$ and $x$ which correspond to the output of the embedding sub-model $\phi(\cdot)$.
The adversarial sample $x_{adv}$ is obtained by solving the optimization problem
$\begin{split}x_{adv}=\arg\min_{x_{adv}}L_{adv}(x_{adv},x,y)\end{split}$ (3)
Optimization is performed using the Adam optimizer with manually tuned learning rates of 1.0 for the Hate and VideoTQ datasets and 20.0 for the ICU dataset.
The detailed procedure for feature selection, perturbation computation, and constraint handling is provided in Appendix A. The full attack design and theoretical framework are described in detail in [8, 9].
@@ 4.5 Regression Models
Of the datasets used in our evaluation, only the ICU dataset contains dependent features whose relationships with other attributes are known to exist but are not explicitly defined or fully observable in the data. Therefore, we applied the regression-based perturbation technique introduced in Section˜3 exclusively to this dataset. Specifically, we identified four dependent features that require regression-based adjustment: ‘apache_3j_bodysystem,’ ‘apache_3j_diagnosis,’ ‘d1_mbp_invasive_max,’ and ‘d1_mbp_invasive_min.’ For each of these features, we trained a separate regression model using the same dataset employed for training the surrogate models.
The regression models were implemented using gradient boosting with 200 estimators and a maximum depth of six. However, different learning rates were required for optimal performance: 0.1 for the ‘apache_3j_bodysystem’ and ‘apache_3j_diagnosis’ features, and 0.01 for the ‘d1_mbp_invasive_max’ and ‘d1_mbp_invasive_min’ features. These hyperparameters were determined through manual tuning to optimize model performance for each feature, with the different learning rates reflecting the varying complexity and scale of the target features. Since satisfactory performance was achieved across all models, no further hyperparameter optimization was conducted.
@@ 4.7 Evaluation Setup
To ensure a robust and meaningful evaluation of adversarial effectiveness, we constructed the attack set by filtering the test set to include only samples that were correctly classified by both the target and surrogate models prior to the attack. This guarantees that any misclassification following the perturbation can be directly attributed to the adversarial manipulation rather than to pre-existing model errors. After filtering, $85\%$, $69\%$, and $91\%$ of the test samples were retained for the Hate, ICU, and VideoTQ datasets, respectively. We then randomly selected a balanced number of samples from each class, resulting in attack sets of 182, 1,000, and 1,000 samples, respectively for the three datasets.
For the purpose of attack evaluation, we assess only adversarial samples that successfully fooled the target model (or both the surrogate and target models in the case of transferability-based attacks). This allows us to analyze the properties of adversarial samples that actually compromise model reliability, as failed attacks do not pose a real threat to downstream users.
When analyzing intrinsic properties of adversarial samples, such as $L_{p}$ distances or anomaly scores, we compare them to the distribution of benign (non-adversarial) training samples. Importantly, this benign reference set includes all training samples, regardless of whether they were correctly classified. This methodological choice captures the full variability of natural data, reflecting a realistic deployment scenario in which even misclassified samples influence the model’s perceived class distribution. As a result, anomaly scores are computed relative to a distribution that detectors might actually encounter in practice.
@@ 4.8 Evaluation Metrics
The main objective of our evaluation is to answer the research questions presented in Section˜1 and analyze the characteristics of and differences between black-box query- and transferability-based adversarial attacks with respect to three key factors: the attacker’s risk (RQ#1), the attacker’s effort (RQ#1), and the attack quality (RQ#2, RQ#3). To perform a thorough analysis, we used the evaluation metrics commonly used in the literature and the metrics proposed in this work.
To assess the attacker’s risk we used three metrics (illustrated in Fig.˜1): i) the percentage of samples that successfully misled the model (attack success rate); ii) the perturbation magnitude computed based on the average number of modified features ($L_{0}$), which reflects practical constraints such as the cost and feasibility of changing specific features in real-world tabular data, and the average Euclidean distance between the original and adversarial samples ($L_{2}$), included to support comparison with prior work and to demonstrate its limitations in capturing detectability in heterogeneous tabular datasets; and iii) the average number of queries required.
We exclude the $L_{\infty}$ norm since, like $L_{2}$, it depends heavily on feature scaling and poorly reflects perturbation detectability or practical impact in tabular domains with mixed feature types and scales [8]. Furthermore, $L_{\infty}$ is not used in prior work on adversarial attacks in tabular domains, reflecting its lack of applicability and relevance in this context.
To assess the attacker’s effort we used three metrics (illustrated in Fig.˜1): i) the time it took to craft an adversarial sample; ii) the number of data samples needed to execute the attack; and iii) the memory resources and computing capabilities required to craft the adversarial sample and perform the optimization process.
The attack quality is reflected in the crafted adversarial sample, which were assessed using three metrics (illustrated in Fig.˜1): i) the anomaly detection rate of adversarial samples compared to benign samples, using IF and AE models, where a higher rate indicates less consistency and coherence (RQ#2); ii) the percentage of samples with at least one feature with an anomalous SHAP value (Importance-Based Anomaly Detection Rate) (RQ#3); and iii) the average anomalous SHAP features per sample (RQ#3).
All three quality metrics are computed using our CSAD approach (see Section˜3.3), which evaluates adversarial deviations relative to the distribution of benign samples from the same predicted class. In this context, lower values across these metrics indicate higher-quality adversarial samples, since they reflect less abnormal behavior within the predicted class. While feature space anomalies (metric i) are model-agnostic and computed once per attack, SHAP-based anomalies (metrics ii and iii) depend on the model’s internal reasoning and are therefore assessed separately for each target model.
@@ 5.1 Models’ Performance
Table˜2 summarizes the performance of all target and surrogate models examined, presenting the accuracy, F1 score, precision, and recall values obtained by each model on the test set.
@@
As can be seen, all models were high-performing. The target models for the Hate dataset obtained accuracy values ranging from $88\%-91\%$], F1 scores ranging from $87\%-92\%$], precision values ranging from $89\%-94\%$], and recall values ranging from $82\%-92\%$]. The target models for the ICU dataset obtained accuracy values ranging from $77\%-80\%$], F1 scores ranging from $75\%-81\%$], precision values ranging from $81\%-84\%$], and recall values ranging from $69\%$-$77\%$]. The target models for the VideoTQ dataset obtained accuracy values ranging from $98\%-99\%$], F1 scores ranging from $99\%-100\%$], precision values ranging from $91\%-97\%$], and recall values ranging from $6\%-99$$9\%$].
@@ 5.2 Attacker’s Risk
Tables˜3 and 4 and Fig.˜2 present the results for the attacker’s risk metrics (see Section˜4.8) for the query- and transferability-based attacks performed against different target models on each dataset.
@@
Table˜3 presents details on the following attack success rates: Target Success Rate (SR): The percentage of adversarial samples that successfully mislead the target model; Surrogate Success Rate (Surrogate SR): The percentage of adversarial samples that successfully mislead the surrogate model; Transfer Success Rate (Transfer SR): The percentage of adversarial samples that successfully transfer from the surrogate model to the target model, calculated as the ratio of the Surrogate SR. Overall Success Rate (Overall SR): The percentage of adversarial samples that successfully mislead both the surrogate and target models, computed as the ratio of the entire attack set.
As detailed in Table˜3, the query-based attacks have consistently high success rates (SR) on all target models across the Hate and ICU datasets. For example, on the Hate and ICU datasets, both attacks achieve success rates ranging between $98\%-100\%$] across all target models. In comparison, the success rate varies more for the transferability-based attacks. The success of these attacks relies on the performance of the surrogate models and the transferability of adversarial samples. For instance, on the Hate dataset, despite successfully fooling surrogate models in $98.9\%$ of cases, the transferability-based attacks demonstrated limited effect on the target models, with success rates falling below $12\%$ when employing random feature selection (Transfer random). However, when using importance-based feature selection (Transfer imp.), the success rates improved, ranging between $20\%-86\%$]. A similar phenomenon can be observed on the ICU dataset, where transferability-based attacks continued to perform variably, depending on the feature selection method employed. On the VideoTQ dataset, unlike the other datasets, transferability-based attacks had more consistent performance, achieving a $\sim$$50\%$ success rate overall. In contrast, the query-based attacks displayed more variability, with success rates ranging between $65.9\%-82.6\%$] for the boundary attack and between $50\%-67.9\%$] for the HopSkipJump attack.
According to pairwise proportions z-tests with Holm correction, the differences in success rates between query- and transferability-based attacks were statistically significant across all datasets ($p<0.05$). Non-significant results were observed only within the transferability group, specifically between variants that achieved similarly moderate success (e.g., GB imp. vs LGB imp. importance-based selection attacks on Hate and ICU, and the different transfer variants on VideoTQ). Full pairwise results are provided in Appendix C.
Table˜4 presents the average number of queries required to generate adversarial samples for query-based attacks. As can be seen, the examined query-based attacks require a significant number of queries, ranging from hundreds to hundreds of thousands. In contrast, transferability-based attacks require just a single query for all datasets.
Fig.˜2 illustrates the distribution of both the number of modified features ($L_{0}$ ) and the Euclidean distance ($L_{2}$) between the adversarial attacks and the original samples.
@@
Statistical validation using pairwise Mann–Whitney U tests with Holm correction confirmed these patterns across datasets. For both $L_{0}$ and $L_{2}$, query-based attacks produced significantly larger perturbations than transferability-based attacks ($p<0.05$), consistent with the visual trends in Fig.˜2. Non-significant differences appeared only within the transferability group, where multiple variants produced similarly small distortions, and on the VideoTQ dataset, where the boundary attack in some cases reached $L_{0}$ and $L_{2}$ values comparable to those of transfer-based attacks. Detailed test results are provided in Appendix C.
@@ 5.3 Attacker’s Effort
To assess the attacker’s effort, we compared the performance of the query- and transferability-based attacks in terms of the computational time, the required amount of data, and the computational resources needed(see Section˜4.8).
With respect to the computational time, the attacker’s objective is to minimize the amount of time it takes to craft an adversarial sample. In query-based attacks, the computational time includes: i) the time it takes to query the target model (denoted as $\alpha$); and ii) the time it takes to perform a single optimization iteration (denoted as $\beta$). The total amount of time it takes to craft an adversarial sample is calculated in Eq.˜4:
@@
where $x$ is a data sample, $M$ is the target model, $n$ is the number of queries required to craft a successful adversarial sample, and $t$ is the time it takes to perform a single query. When considering solely the query-related aspects of the required time, query-based attacks are at a disadvantage compared to transferability-based attacks, since many more queries to the target model are required (see Section˜5.2). However, in transferability-based attacks, there is another aspect to consider – the surrogate model’s training and querying time. The total amount of time it takes to craft an adversarial sample is calculated in Eq.˜5, where we assume that the time it takes to perform a query to the target model is identical to the time required to perform a query to the surrogate model.
@@
In our evaluation, we found that the total amount of time it took to produce a single adversarial sample in query-based attacks was, on average, $518.4$ seconds. However, in transferability-based attacks, it was $\sim$$243.5673$ seconds: $\sim$$240.49$ seconds to train the surrogate and feature importance models, $0.0773$ seconds to generate the adversarial sample using the surrogate model, and $\sim$33 seconds to query the target model. Based on this, we can conclude that although transferability-based attacks require training additional models, the total computational time required to craft a single adversarial sample is shorter than in query-based attacks.
@@
In conclusion, while query-based attacks require less data, transferability-based attacks require less computational time and resources. However, they are limited by lower success rates compared to query-based attacks, as discussed in Section˜5.2.
@@ 5.4 Attack Quality
In our evaluation, we assessed query- and transferability-based attacks in terms of the quality of the adversarial samples produced. This quality can be examined by considering both the adversarial sample’s coherence and its impact on the target model’s decision-making process. To measure the coherence of adversarial samples, we propose using metrics based on AE and IF models. To evaluate the impact on the target model’s decision-making process, we introduced two new metrics derived from the SHAP framework (see Section˜4.8).
Feature Space Coherency.
@@
Fig.˜3 presents the anomaly detection rates obtained by applying IF anomaly scores and AE reconstruction errors to adversarial samples across different attacks, target models, and datasets. Each detection rate represents the true positive rate (TPR), indicating the percentage of adversarial samples identified as anomalous. For the AE, a sample is considered anomalous if its reconstruction error exceeds a predefined threshold (see  Eq.˜1). The resulting false positive rate (FPR) for benign test samples is small and fixed within a dataset: 0.73% and 4.99% for class 0 and class 1 on the Hate dataset, 3.06% and 3.58% on the ICU dataset, and 0.07% and 1.45% on the VideoTQ dataset. For comparability, IF was calibrated to the same FPR levels per dataset and class. Given that the FPR values are consistently very low (ranging between 0.07% and 4.99% across all datasets), the reported detection rates effectively represent the performance of each method; the false negative rate (FNR) can be directly obtained as $1-\mathrm{TPR}$.
(a) Hate — AE
(b) Hate — IF
(c) ICU — AE
(d) ICU — IF
(e) VideoTQ — AE
(f) VideoTQ — IF
Figure 3: Attack quality evaluated based on the feature space anomaly detection. Bars show the percentage of adversarial samples flagged as anomalies (TPR; higher is better) by an AE and an IF for each attack and target model, across datasets. Thresholds are fixed per dataset and class on a benign validation split; the resulting benign FPR is less than 5% across all datasets and classes.
[…]
On the VideoTQ dataset, query-based attacks achieved only low anomaly detection rates, with boundary attacks reaching at most $\approx 20$$25\%$ and HopSkipJump slightly lower, where IF achieved better detection rates than AE. In contrast, transferability-based attacks were essentially undetected by both detectors, with detection rates close to $0\%$ across all variants.
According to proportions z–tests with Holm correction, the observed differences in anomaly detection rates were statistically significant, with IF detecting query-based attacks at substantially higher rates than transferability-based attacks across all datasets ($p<0.001$), with only isolated exceptions, such as the random-based selection attack on the ICU dataset. A similar trend was observed for AE: detection rates for query-based attacks were consistently and significantly higher than for transferability-based attacks ($p<0.001$), despite AE achieving overall lower detection levels than IF, particularly on boundary and HopSkipJump attacks. Full pairwise results for both detectors are provided in Appendix C.
@@
As shown in Fig.˜3, the two query-based attacks demonstrate distinct behaviors in terms of the anomaly detection rate. Most adversarial samples generated by the boundary attack were detected as anomlies, particularly by IF. In contrast, adversarial samples from the HopSkipJump attack proved much harder to detect by both models, closer in magnitude to those observed for transferability-based attacks.
These results reveal two consistent patterns. First, query-based attacks, particularly the boundary attack, are substantially easier to detect, with IF outperforming AE. Transferability-based attacks, by contrast, remain less detectable by both models across all datasets, with only isolated cases where AE achieves a small advantage. Second, the comparison between AE and IF indicates that IF is systematically more effective in capturing adversarial samples that are “far” or structurally shifted in the feature space, as reflected in its dominance on boundary and random-based selection attacks. AE, in contrast, showed only sporadic and inconsistent advantages, with its effectiveness largely limited outside boundary attack scenarios. The substantial gap between query- and transferability-based anomaly detection underscores the persistent challenge posed by transferability-based gradient attacks, particularly in realistic settings where these attacks are almost entirely undetectable as anomalies.
@@
Notably, despite HopSkipJump and transferability-based attacks often producing adversarial samples with lower $L_{2}$ distances (as detailed in Section˜5.2), these samples were still identifiable as anomalous in some cases. This suggests that even subtle perturbations introduced by these attacks can be identified as anomalous.
@@
Figs.˜4, 5 and 6 presents the distribution of SHAP values for adversarial versus benign samples across different attacks and datasets for each target model. To simplify the presentation, we focus on the top-four most important features, selected based on their average SHAP values computed on the benign training dataset. As shown in the figures, for most target models and datasets, there is at least one feature where the SHAP value distribution across attacks varies significantly. Analyzing the SHAP-based feature importance of a queried sample can therefore be an effective technique for identifying attacks.
Table˜5 presents the results of our evaluation of attack quality using SHAP-based metrics. For each attack and target model, we report the Importance-Based Anomaly Detection Rate (percentage of samples with at least one anomalous feature based on SHAP values) and the Average Anomalous SHAP Features per Sample (as described in Section˜4.8). Lower values indicate that there is less impact on the model’s decision-making process, meaning that adversarial samples were more consistent with benign samples in terms of feature importance (i.e., the model relies on similar features for prediction).
@@
On the ICU dataset, both query- and transferability-based attacks generally produced less anomalous SHAP values for most target models (LGB, XGB, and RF), with importance-based anomaly detection rates typically below 7% and very low values of average anomalous SHAP features per sample ranging from 0.0 to 0.1. However, when targeting the GB model, transferability-based attacks resulted in notably lower importance-based anomaly detection rates (1.8% to 10.8%) compared to query-based attacks (29.0% and 19.4% for the boundary and HopSkipJump attacks, respectively). In this case, the average number of anomalous SHAP features per sample was slightly higher for query-based attacks (0.3 to 0.4) than for transferability-based attacks (0.0 to 0.1), although still relatively low overall.
@@
According to Mann–Whitney U tests (for the average anomalous SHAP features per sample) and proportions z–tests (for the importance-based anomaly rate), both with Holm correction, the observed differences were statistically significant across all datasets ($p<0.001$). On the Hate and ICU datasets, query-based attacks yielded substantially higher importance-based anomaly rates than transferability-based attacks, with medium effect sizes. On VideoTQ, query-based attacks also showed higher anomaly rates overall, although the differences in the average anomalous SHAP features per sample were smaller, with significance observed across all target models except under the random-based selection attack when targeting the GB model. Within the transferability-based group, omnibus tests revealed heterogeneity among importance-based selection variants, but no single strategy consistently outperformed the others. Among the query-based methods, the boundary attack generally produced more anomalous SHAP values than HopSkipJump, a difference that reached statistical significance in several cases. Full test results are provided in Appendix C.